"A line passes through the points listed in the table below. Examine the ordered pairs in the table. What relationship do you notice between the x and y values? Express this relationship by writing an aquation that every point in the table must satisfy."
x | y
1 | 5
2 | 4
3 | 3
4 | 2
5 | 1
I know it's a linear equation and there is a different of 1 between each value. I just can't seem to put it in equation form for some reason.

Answer by Edwin McCravy(20060)

x | y
1 | 5 <-- notice that 1+5=6
2 | 4 <-- notice that 2+4=6
3 | 3 <-- notice that 3+3=6
4 | 2 <-- notice that 4+2=6
5 | 1 <-- notice that 5+1=6
So what you were to notice was that in every case
above,
x + y = 6
Edwin